### v1

# journal: TOG

Abstract.
We introduce MeshPad, a generative approach that creates 3D meshes from sketch inputs. Building on recent advances in artistic-designed triangle mesh generation, our approach addresses the need for interactive artistic mesh creation. To this end, we focus on enabling consistent edits by decomposing editing into ‘deletion’ of regions of a mesh, followed by ‘addition’ of new mesh geometry. Both operations are invoked by simple user edits of a sketch image, facilitating an iterative content creation process and enabling the construction of complex 3D meshes. Our approach is based on a triangle sequence-based mesh representation, exploiting a large Transformer model for mesh triangle addition and deletion. In order to perform edits interactively, we introduce a vertex-aligned speculative prediction strategy on top of our additive mesh generator. This speculator predicts multiple output tokens corresponding to a vertex, thus significantly reducing the computational cost of inference and accelerating the editing process, making it possible to execute each editing step in only a few seconds. Comprehensive experiments demonstrate that MeshPad outperforms state-of-the-art sketch-conditioned mesh generation methods, achieving more than $22\%$ mesh quality improvement in Chamfer distance, and being preferred by $90\%$ of participants in perceptual evaluations.
journal: TOG
[…]
Haoxuan Li
Ziya Erkoc
Lei Li
Daniele Sirigatti
Vladislav Rosov
Angela Dai
Matthias Nießner
Figure 1. MeshPad enables interactive mesh creation and editing with sketches. We decompose this complex task into two sketch-conditioned operations: addition and deletion. Top: our method allows for a user to create and modify artistic-designed triangle meshes by simply drawing and editing 2D sketches, achieving intuitive and interactive 3D modeling. Bottom (left): our method generates a variety of complex yet compact meshes. Bottom (right): our interactive user interface allows users to iteratively edit the mesh, with each edit step taking a few seconds.
1. Introduction
Figure 2. Method Overview. We show the decomposition of mesh creation and editing into addition and deletion operations (left and right, respectively). ADD: in mesh addition, a transformer generates new mesh regions corresponding to newly added strokes (in red) in the input sketch. The generated mesh triangles are then merged with the existing mesh. DEL: in mesh deletion, we show a deletion operation applied to the mesh addition output. We erase sketch strokes to remove corresponding mesh regions, and overlay the erased strokes in red alongside untouched regions to provide more context to the deletion network. We then predict which mesh vertices correspond to regions to be deleted, and prune the mesh accordingly. After each addition or deletion operation, we automatically generate an updated sketch corresponding to the current output mesh, to enable further sketch-based editing.
Triangle meshes are one of the most predominant 3D representations used in 3D production applications, from video games to virtual reality and movies. Mesh creation and editing is thus a central element of computer graphics. In contrast to 3D representations such as voxels (Dai et al., 2017, 2020; Ren et al., 2024), points (Vahdat et al., 2022; Zhou et al., 2021; Yang et al., 2019), or neural implicit representations (Hertz et al., 2022; Mildenhall et al., 2021; Kerbl et al., 2023), triangle meshes represent surfaces in a compact, structured fashion, as well as enabling efficient fine-scale detail and naturally integrating into modern rendering and editing pipelines, achieving high fidelity with relatively few primitives.
Recent advances in generative 3D models have shown significant potential in generating 3D meshes directly (Siddiqui et al., 2023; Chen et al., 2024b, e; Tang et al., 2024; Hao et al., 2024); however, these output meshes are not editable, which is a crucial component for artistic design in content creation. More specifically, artistic content creation is an iterative process that encompasses not only the initial generation, but also requires multiple cycles of manipulation and editing in order to refine an output to achieve a precise artistic vision. In order for such editing to fit with content creation pipelines, various edits must be performed interactively, affecting only the intended region of the mesh.
We thus propose MeshPad, an interactive sketch-based approach for 3D mesh generation and editing. From an input sketch of a shape drawn by a user, we produce a corresponding 3D mesh. The resulting mesh can then be edited simply by editing the sketch. To achieve efficient and precise mesh editing, we decompose this into simpler subtasks: deletion and addition of mesh geometry by removing or adding strokes in the sketch. Crucially, this modeling paradigm also leads to easy supervised training of both deletion and addition of elements in the mesh (by simply removing and adding back parts of ground-truth meshes), without requiring collection of real mesh editing sequences for supervision. Our editing-based approach allows for the generation of more complex triangle meshes by iteratively applying a series of edits to construct the final shape.
Algorithmically, MeshPad leverages a hybrid approach combining an autoregressive network for addition and a token-classification network for deletion. Unlike existing artistic-designed mesh generation methods, our addition network generates partial shapes only corresponding to new sketch strokes, rather than repeatedly synthesizing the entire mesh. This naturally solves the problem of preserving unedited mesh regions during editing. Additionally, based on the triangle sequence representation (Chen et al., 2024e), we introduce a vertex-aligned speculative decoder to accelerate autoregressive generation and reduce computational time. Instead of predicting one vertex coordinate at a time, the speculator allows us to predict $3$ coordinates, i.e., one vertex, at a time. We find that jointly training the speculative head and aligning it with the vertex tokens reduces generation time while maintaining generation quality. Furthermore, we implemented a user interface that allows users to interactively edit 3D meshes by drawing 2D sketches. Extensive experiments show that our method outperforms state-of-the-art methods in sketch-conditioned mesh generation, producing cleaner meshes that more accurately align with the sketch input and excelling in partial mesh editing tasks. Finally, our speculative prediction acceleration enables interactive mesh generation and editing within seconds while maintaining overall shape quality.
In summary, our contributions are:
  • We introduce a novel method for interactive mesh creation and editing by decomposing the process into addition and deletion. This enables easy training without requiring collection of edited ground-truth meshes, and enables finer-grained control over an iterative 3D mesh creation process.
  • Our vertex-aligned speculative classification head for addition notably accelerates the triangle mesh generation by $\times2$ without quality loss.
2. Related Work
3D Mesh Generation. Direct mesh generation offers the compelling advantage of producing outputs that closely resemble artist-crafted 3D shapes. Early approaches proposed various parameterizations of an irregular mesh structure, including altases (Groueix et al., 2018) and graphs (Dai and Nießner, 2019). Recently, Polygen (Nash et al., 2020) and MeshGPT (Siddiqui et al., 2023) have demonstrated the remarkable potential of transformers in autoregressively generating artistic meshes. Follow-ups such as MeshAnything (Chen et al., 2024b), MeshAnything V2 (Chen et al., 2024e), EdgeRunner (Tang et al., 2024), and Meshtron (Hao et al., 2024) each proposed improvements in tokenization, attention mechanisms, or shape coverage, pushing the limit of direct mesh generation. PolyDiff (Alliegro et al., 2023) instead uses a diffusion backbone for polygonal mesh synthesis. Despite solid progress, these methods focus on full-shape generation and cannot easily handle local editing tasks.
Conditional 3D Generation. The history of sketch-driven methods starts with techniques that inflate 2D contours (Igarashi et al., 1999; Nealen et al., 2007) or retrieve 3D models from sketch-based queries (Eitz et al., 2012; Funkhouser et al., 2003). Modern sketch-based 3D generation methods typically rely on either single-shot generation SDF diffusion models (Zheng et al., 2023; Robin Borth, 2024) or direct editing of neural fields (Mikaeili et al., 2023), which usually struggle to output tidy, artistically crafted meshes.
With the advancement in large text-to-image models, DreamFusion (Poole et al., 2023) proposes score distillation sampling with 2D diffusion for text-to-3D generation, followed by Magic3D (Lin et al., 2023), Fantasia3D (Chen et al., 2023b), and Meta 3D AssetGen (Siddiqui et al., 2024). Text conditions lack precision for fine-grained control (Xu et al., 2023) and additional image condition is often required (Qian et al., 2024; Bala et al., 2024). In contrast, our method uses iterative sketch conditioning for fine-grained shape control.
3D Shape Editing. With the same concept of SDS, many 3D shape editing methods rely on neural representations (Khalid et al., 2025; Liu et al., 2025), and use region-specific masks (Chen et al., 2024a; Barda et al., 2024; Chen et al., 2024d). SPAGHETTI (Hertz et al., 2022), SALAD (Koo et al., 2023), and PartGen (Chen et al., 2024c) advance part-level neural shape editing by enabling disentangled control, diffusion-based manipulation, and multi-view diffusion for generating and editing meaningful 3D parts, respectively.
Recently, sketch-based 3D editing has emerged as an intuitive alternative to text- and image-based methods. SKED (Mikaeili et al., 2023) and SketchDream (Liu et al., 2024) leverage sketch-conditioned 2D diffusion models for 3D modeling. Other approaches, including SENS (Binninger et al., 2024), Doodle your 3D (Bandyopadhyay et al., 2024), and Masked LRM (Gao et al., 2024), focus on abstraction-aware part manipulation, robust part isolation, and masked reconstruction for fast local edits, respectively.
However, such works struggle to produce an artistic mesh. Techniques based on polygonal meshes or other explicit representations can rely on an expensive optimization (Aigerman et al., 2022; Gao et al., 2023), deformation priors (Tang et al., 2022), specialized parametric families (Elrefaie et al., 2024) or procedural generators (Zhao et al., 2024). Such approaches do not generalize easily to arbitrary geometry and have difficulties making local edits without conflicting with the global constraints. By contrast, our approach addresses these issues using a sequence-based mesh representation to perform partial editing by removing and then adding triangles.
Speculative Decoding. Finally, to reduce inference time for our autoregressive network, we adopt a concept of “speculative decoding”, analogous to language Transformers that generate multiple tokens in one pass (Leviathan et al., 2023; Chen et al., 2023a; Miao et al., 2023; Wertheimer et al., 2024). We design a vertex-aligned speculator for the mesh sequence to accelerate partial additions, enabling agile and expressive iterative mesh editing.
3. Method
MeshPad enables interactive mesh creation and editing by iteratively performing addition and deletion operations guided by input sketches. An input sketch image $\mathcal{I}$ guides the generation of a 3D mesh $\mathcal{M}$ as a sequence of mesh triangles. $\mathcal{M}$ can be further edited by removing strokes from $\mathcal{I}$ to delete mesh regions, or adding new strokes to introduce new geometry. To achieve interactive rates for editing, we introduce a vertex-aligned speculator into our mesh addition transformer, significantly accelerating its autoregressive generation. An overview of our approach is shown in Fig. 2.
Figure 3. Model Architecture. We use the Open Pre-trained Transformer (OPT) as the Transformer backbone to process the embedded triangle sequence and the sketch embeddings from a pre-trained image foundation model, RADIO. Top (right): the deletion network is a classification network that labels each input mesh vertex for removal. The deletion head takes the $xyz$ coordinate embeddings of a vertex as input and produces a masking label for the vertex. Bottom (right): the addition network is an autoregressive generation model that predicts new mesh tokens with a speculator. During the speculative prediction, we align the speculator to the vertex axis so that it always takes the $x$ coordinate as input and predicts the vertex’s $y$ and $z$ coordinates. Note that E, M, V, and ¡s¿ represent Transformer hidden states, vertex masks, vertex coordinate tokens, and the ¡split¿ token, respectively.
3.1. Mesh Creation and Editing
We decouple sketch-based mesh creation and editing into two subtasks: addition and deletion. Both mesh operations are conditioned on a sketch image $\mathcal{I}$, a bitmap where colored line strokes indicate regions to be added or deleted. We define a mesh $\mathcal{M}$ as a sequence of triangles ${}\mathcal{M}=\{\mathcal{F}\}$, where =$\mathcal{F}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}\}$ and $\boldsymbol{v}\in\mathbb{R}^{3}$ represents 3D vertex locations.
Line strokes in the input sketch image $\mathcal{I}$ are divided into two mutually exclusive sets, $\mathcal{I}_{k}$ and $\mathcal{I}_{r}$, colored black and red in our visualizations, respectively. When performing an addition or deletion, $\mathcal{I}_{r}$ corresponds to the user edit, i.e., the part to be added or deleted, respectively, while $\mathcal{I}_{k}$ represents the untouched sketch regions. We denote the mesh parts corresponding to $\mathcal{I}_{r}$ and $\mathcal{I}_{k}$ as mutually exclusive sets $\mathcal{M}_{r}$ and $\mathcal{M}_{k}$, respectively, and thus $\mathcal{M}_{r}\cup\mathcal{M}_{k}=\mathcal{M}$. The operation–addition or deletion–is determined by whether the $\mathcal{I}_{r}$ is added as new strokes (addition), or erased from existing strokes (deletion). This also enables simple supervision of both deletion and addition networks by removing or adding back random regions of mesh geometry of 3D shapes, without requiring ground-truth sequences of mesh edits (see Sec. 3.4 for more detail).
Sketch-conditioned Mesh Deletion. Given a current mesh state $\mathcal{M}$ and a sketch image $\mathcal{I}$ containing a non-empty set of deletion strokes $\mathcal{I}_{r}$, the goal is to obtain a mesh $\mathcal{M}_{k}\subsetneq\mathcal{M}$ that excludes parts corresponding to $\mathcal{I}_{r}$. This is formulated as a binary classification task over the sequence of triangles composing $\mathcal{M}$; those predicted as corresponding to $\mathcal{I}_{r}$ are then removed.
We employ a transformer-based architecture (Sec. 3.2) to predict a binary label for each vertex in the mesh sequence, forming a set of vertices $\mathcal{V}^{\prime}_{r}$ corresponding to $\mathcal{I}_{r}$. Then we predict the target mesh $\mathcal{M}^{\prime}_{k}$ by deleting triangles with any vertex predicted as deleted:
(1) $\mathcal{M}^{\prime}_{r}=\left\{\mathcal{F}\in\mathcal{M}|\exists\boldsymbol{v}\in\mathcal{F}:\boldsymbol{v}\in\mathcal{V}^{\prime}_{r}\right\};\quad\mathcal{M}^{\prime}_{k}=\mathcal{M}\setminus\mathcal{M}^{\prime}_{r}.$
We train this by automatically generating 2D sketches and removal regions from 3D shapes in a self-supervised fashion (Sec. 3.4) to obtain the ground truth $\mathcal{V}_{r}=\bigcup_{\mathcal{F}\in\mathcal{M}_{r}}\mathcal{F}$ and set the binary label of each vertex as:
(2) $l_{\boldsymbol{v}}=\begin{cases}0&\text{if }\boldsymbol{v}\in\mathcal{V}_{r},\\ 1&\text{otherwise}.\end{cases}$
This deletion is trained with a binary cross-entropy loss between the predicted binary probability per vertex and the vertex label. Once deletion is performed, we assign ${\mathcal{M}\leftarrow\mathcal{M}^{\prime}_{k}}$ and ${\mathcal{M}_{k}\leftarrow\mathcal{M}^{\prime}_{k}}$, and automatically generate the updated $\mathcal{I}_{k}$ corresponding to the new output mesh $\mathcal{M}$ (Sec. 3.4), which can be further edited.
Sketch-conditioned Mesh Addition. The addition task is the inverse process of deletion. The input is a mesh $\mathcal{M}_{k}$ and a sketch $\mathcal{I}$ containing non-empty $\mathcal{I}_{r}$, and the goal is to generate the mesh $\mathcal{M}$. We use a similar architecture as with deletion to autoregressively generate tokens of $\mathcal{M}^{\prime}_{r}$ and merge it with $\mathcal{M}_{k}$ to obtain the predicted mesh $\mathcal{M}^{\prime}$. This is supervised by a cross-entropy loss between the ground-truth mesh tokens of $\mathcal{M}_{r}$ and the predicted token probabilities. After merging, we assign ${\mathcal{M}\leftarrow\mathcal{M}^{\prime}}$ and ${\mathcal{M}_{k}\leftarrow\mathcal{M}^{\prime}}$.
We automatically generate an updated sketch $\mathcal{I}_{k}$ reflecting the latest mesh geometry. This enables iterative mesh generation and editing. A user first draws a sketch to produce an initial mesh or loads in a mesh from which we automatically generate a corresponding sketch. Editing can then be performed on this initial mesh and sketch; a user can erase a sketch part with an eraser tool, labeling part of $\mathcal{I}_{k}$ as $\mathcal{I}_{r}$ and triggering deletion of the corresponding mesh part. The user can also draw new strokes alongside the current $\mathcal{I}_{k}$ to introduce new mesh geometry. This iterative editing process avoids re-synthesis of the whole shape, preserving mesh structures in unedited regions. Note that mesh generation from scratch is treated as addition to an empty mesh.
3.2. Network Architecture
The detailed model structure is shown in Fig. 3. We use a pre-trained image encoder RADIO 2.5-h (Ranzinger et al., 2024) to encode the sketch input as tokens. We follow MeshAnythingV2 (Chen et al., 2024e) to use the same Open Pre-trained Transformer (OPT) (Zhang et al., 2022) network as our backbone. We also adopt the tokenizer $\mathcal{T}$ from MeshAnythingV2 to transform a mesh into a sequence $S=\mathcal{T}(\mathcal{M})$. The sequence $S$ is defined as an ordered list consisting of control tokens (¡split¿, ¡start¿, ¡end¿) and vertex coordinate tokens $V$. The ¡start¿ and ¡end¿ tokens indicate the bounds of the sequence. The ¡split¿ token splits the sequence into subsequences containing only vertex tokens, which define a series of adjacent triangles (Chen et al., 2024e).
To perform mesh addition, we use the OPT model to autoregressively predict the next sequence token $S_{r}^{\prime(i+1)}$ given the sketch $\mathcal{I}$, input mesh sequence $S_{k}=\mathcal{T}(\mathcal{M}_{k})$, and the previously generated tokens $S_{r}^{\prime(1\ldots i)}$:
(3) $P\left(S_{r}^{\prime(i+1)}|S_{k},\mathcal{I},S_{r}^{\prime(1\ldots i)}\right)=\text{OPT}\left(S_{k},\mathcal{I},S_{r}^{\prime(1\ldots i)}\right).$
After generating the sequence $S^{\prime}_{r}$, we obtain an output mesh by detokenizing $S^{\prime}_{r}$ and merging it with the input partial mesh:
(4) $\mathcal{M}^{\prime}=\mathcal{M}_{k}\cup\mathcal{T}^{-1}(S^{\prime}_{r}).$
Instead of generating one token at a time, we introduce a vertex-aligned speculator to predict multiple tokens in a single run (Sec. 3.3).
The deletion network predicts a mask label for each vertex with a deletion head, which processes the encoded states of every three coordinate tokens corresponding to a vertex. While using the same OPT architecture, the attention layers are switched to perform bi-directional attention for each position to capture the global context. After the inference, we aggregate $\mathcal{V}^{\prime}_{r}$ according to Eq. 2, and then obtain the deleted mesh with Eq. 1.
3.3. Vertex-aligned Speculator
As we perform mesh addition autoregressively, generating an $n$-token sequence requires $n$ forward passes, significantly limiting mesh generation speed for interactive applications. One effective way of accelerating generation is to use a speculator to generate multiple tokens in a single run. In contrast to natural language, a tokenized mesh sequence contains strong low-level structure – each vertex $\boldsymbol{v}$ is defined by exactly $3$ tokens $V_{\{x,y,z\}}$ representing its $x$, $y$, and $z$ coordinates. Aligning the speculator with vertex tokens by letting it predict only $V_{\{y,z\}}$, as shown in Fig. 3, ensures a consistent context of input and output for the speculator.
We adopt the MLP speculator described in (Wertheimer et al., 2024). In natural language processing tasks, a speculator can be trained with a frozen pre-trained large language model to accelerate generation. However, in our experiments, we found that this results in performance degradation. We thus train our speculator jointly with the the mesh addition task, so that the addition transformer learns to inform the speculator with contextual information in hidden states.
As shown in Fig. 3, the speculator predicts the $y$, $z$ coordinates of a vertex $\boldsymbol{v}^{\prime}$ by:
(5) $P\left(V^{\prime}_{\{y,z\}}\right)=\text{Speculator}\left(E_{x},V^{\prime}_{x}\right),$
where $E_{x}$ represents Transformer hidden states corresponding to $V_{x}$. The speculator is trained with a cross-entropy loss between ground truth $V_{\{y,z\}}$ and the prediction. When jointly trained with OPT, the OPT loss function supervises $V_{x}$ and control tokens only.
3.4. Self-Supervised Data Generation for Mesh Editing
By decomposing sketch-conditioned mesh editing into separate deletion and addition subtasks, we can generate training data for each subtask from 3D shape datasets without requiring real mesh editing sequences.
Given a complete 3D mesh sample $\mathcal{M}_{c}$, we select its triangles to form two mutually exclusive subsets $\mathcal{M}_{r}$ and $\mathcal{M}_{k}$, representing the edited and unedited parts, respectively. Note that $\mathcal{M}_{r}\cup\mathcal{M}_{k}=\mathcal{M}_{c}$ is not required, as intermediate meshes during editing can be incomplete before deletion or after addition. To achieve this, we sample two volumes $\mathcal{L}$ and $\mathcal{L}_{k}\subsetneq\mathcal{L}$ within the bounding volume of $\mathcal{M}_{c}$. An example is shown in Fig. 4. For each volume, we crop a partial mesh by selecting all triangles containing at least one vertex within the given volume:
(6) =$\mathcal{M}=\left\{\mathcal{F}\in\mathcal{M}_{c}|\exists\boldsymbol{p}\in\mathcal{F}:\boldsymbol{p}\in\mathcal{L}\right\},$
and the same for $\mathcal{M}_{k}$, $\mathcal{L}_{k}$. We then compute $\mathcal{M}_{r}=\mathcal{M}\setminus\mathcal{M}_{k}$, which is mutually exclusive from $\mathcal{M}_{k}$. We provide more details about the volume sampling and mesh cropping in the supplementary. The obtained meshes are then used to supervise both the addition and the deletion network in the tasks defined in Sec. 3.1.
Figure 4. An example of training data generation: the volume $\mathcal{L}$ is sampled to cover a large portion from the top, while $\mathcal{L}_{k}$ covers the chair backrest. The mesh $\mathcal{M}$ and $\mathcal{M}_{k}$ contains triangles with any vertex in $\mathcal{L}$ and $\mathcal{L}_{k}$, respectively. $\mathcal{M}_{r}$ is the difference between $\mathcal{M}$ and $\mathcal{M}_{k}$. The red and black line strokes in sketch $\mathcal{I}$ corresponds to $\mathcal{M}_{k}$, and $\mathcal{M}_{r}$, respectively.
Sketch Generation. We generate synthetic sketches corresponding to the processed mesh data. During training, this is done automatically through Canny edge detection; at test time, any human-drawn sketch can be used. Fig. 4 shows our data generation process.
Automatic sketch generation is done by applying Canny edge detection on the rendered depth and normal images of $\mathcal{M}$. We then merge the two edge detection results to obtain the corresponding synthetic sketch used for training. To get mutually exclusive sets of line strokes $\mathcal{I}_{\{r,k\}}$, which correspond to different parts of the mesh, we further render a visibility mask of $\mathcal{M}_{r}$ in the same camera view and collect all line strokes within the mask as $\mathcal{I}_{r}$. $\mathcal{I}_{k}$ is then defined as the line strokes not in $\mathcal{I}_{r}$. During training, we randomly sample camera views from the upper unit viewing hemisphere, with azimuth angles in $[-90^{\circ},90^{\circ}]$ and elevation angles in $[0^{\circ},60^{\circ}]$.
4. Experiments
4.1. Implementation Details
Data. Following MeshGPT (Siddiqui et al., 2023) and MeshAnythingV2 (Chen et al., 2024e), we preprocess and filter the ShapeNet dataset (Chang et al., 2015) to obtain approximately $28$k meshes from $55$ categories, which each have a face count $<768$ for training. We sample around $500$ meshes for validation and $1000$ meshes for testing. Other meshes are used for training.
Training. Both the mesh addition and deletion networks are initialized using the OPT weights from the MeshAnythingV2 checkpoint, with the RADIO image model frozen throughout training. We employ a learning rate schedule from $1\times 10^{-4}$ to $1\times 10^{-6}$. For the addition model, training begins with a batch size of $56$ for $128$ epochs, followed by $20$ epochs with a batch size of $224$. The deletion model is trained with a batch size of $32$ for $128$ epochs. Using 4 NVIDIA A100 GPUs, the training process takes approximately $5$ days for the addition network and $2$ days for the deletion network.
User Interface. We develop a Gradio (Abid et al., 2019) user interface to demonstrate our method in an interactive editing environment, accessible from browsers on all types of devices (PC, iPad, etc.). As visualized in Fig. 1, the interface contains a sketchpad for users to draw (addition) and erase (deletion) line strokes and a three.js (Cabello and Contributors, 2010) viewer displaying the mesh in real-time. Users can submit their sketches to the addition or deletion job and can always re-sketch and re-submit the job when the outcome should be iterated on. When an edit is accepted, the interface refreshes the sketch in the sketchpad to match the current mesh for further edit, which corresponds to the process depicted in Fig. 2. We refer to our supplementary video for live demos.
4.2. Baselines and Metrics
In our experiments, we evaluate the generated mesh quality as well as its consistency with the input sketch. For mesh quality, we report the Chamfer Distance (CD) to the target mesh. A smaller CD indicates better geometric consistency with the target. Additionally, for perceptual mesh quality evaluation we employ shading-image-based Fréchet Inception distance (FID) following MeshGPT (Siddiqui et al., 2023) and SENS (Binninger et al., 2024). We also evaluate sketch-based CLIP (Radford et al., 2021) similarity, following LAS (Zheng et al., 2023), to measure the similarity between the generated mesh and the sketch. We additionally introduce sketch-based LPIPS (Zhang et al., 2018) to evaluate sketch-to-mesh correspondence: while CLIP shows the similarity of the global context, LPIPS focuses more on local features. For our method, we analyze runtime with the token generation speed in $tokens/second$ (T/s).
We further conduct both unary and binary perceptual studies. In the unary study, participants evaluate mesh quality (GQ) and sketch matching (GM) for mesh generation. For mesh editing, they rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part. Each is rated on a scale of $1$ to $5$ ($=$best). In the binary study, we ask participants to compare our method with baselines and select their preferred method. We report the percentages of preferences for our method compared to the baselines.
We compare with two state-of-the-art sketch-conditioned 3D generation methods, LAS (Zheng et al., 2023) and SENS (Binninger et al., 2024) as baselines. Since these methods produce signed distance fields or occupancy grids, we use marching cubes to convert their results to meshes. We alternatively use MeshAnythingV2 (Chen et al., 2024e) to post-process LAS and SENS results (LAS-MA and SENS-MA, respectively), producing more artistically-styled meshes for these baselines.
4.3. Experimental Setup
We use the test set described in Sec. 4.1 and automatically generate sketches from one of five views: top-left, top-front-left, top-front, top-front-right, or top-right, for the generation task. To further show the robustness of our method, we use the IKEA dataset (Lim et al., 2013; Sun et al., 2018), which contains $188$ furniture models. Additionally, we randomly select $50$ shapes of airplanes, chairs, and lamps (on which SENS is trained) from the two datasets and manually edit the automatically generated sketches to build an evaluation dataset for sketch-based mesh editing.
As SENS is category-specific and is only trained on airplanes, chairs, and lamps, we evaluate it only on these categories. For the editing task, we let LAS generate from scratch conditioned on the edited sketch since it does not support localized editing; as SENS was designed for editing, and we ran its editing using its original design. For our method, while providing only the original sketch and the edited sketch, we first automatically generate input for deletion by identifying the difference between the two sketches, then perform the deletion-addition operation.
4.4. Comparison to State of the Art
Mesh Generation. Tab. 1 evaluates MeshPad in comparison with state of the art on mesh generation. Quantitatively, our method outperforms baselines in all evaluation metrics, indicating superior mesh quality and sketch-to-mesh correspondence. We also conduct a perceptual study with $35$ participants, and show results in Tabs. 2 and 3, further demonstrating user preference for our generated meshes. Fig. 7 provides qualitative comparison across the methods. While MeshAnythingV2 can convert marching cube meshes to meshes with more artist-reminiscent design, it is not robust enough to handle various artifacts in outputs of LAS and SENS (e.g., floaters or incomplete parts), resulting in an overall worse quality than our method, which bridges directly sketch to mesh.
Table 1. Evaluation on sketch-conditioned mesh generation. We compare with LAS (Zheng et al., 2023), SENS (Binninger et al., 2024), and further use MeshAnythingV2 (MA) (Chen et al., 2024e) to post-process LAS and SENS outputs. The unit of CD is 0.001. MeshPad outperforms baselines in both mesh quality and consistency with input sketches.
[…]
ShapeNet LAS 22.46 0.3309 93.01 47.07
LAS-MA 29.69 0.3595 92.12 20.85
SENS 8.95 0.2753 93.36 81.88
SENS-MA 29.43 0.3348 91.88 42.93
Ours 6.20 0.1790 95.85 9.38
IKEA LAS 20.69 0.3281 93.70 68.51
LAS-MA 26.54 0.3576 93.53 50.75
SENS 8.76 0.2722 94.38 115.33
SENS-MA 19.18 0.3037 94.37 99.08
Ours 6.78 0.1837 96.67 29.67
Table 2. Sketch-based metrics and unary perceptual study ratings (ranged $1$-$5$) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS (Zheng et al., 2023) and SENS (Binninger et al., 2024), with outputs further refined using MeshAnythingV2 (MA) (Chen et al., 2024e). For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.
Sketch Matching Gen. Rating Edit Rating
[…]
LAS 0.3449 92.14 3.2 3.0 2.7 2.2 2.5
LAS-MA 0.3789 91.99 2.8 2.6 2.4 2.0 2.0
SENS 0.3179 91.49 3.4 3.5 2.9 1.8 3.7
SENS-MA 0.3651 90.20 2.7 2.8 2.1 1.6 2.5
Ours 0.2218 95.71 4.3 4.3 4.3 4.2 4.3
Table 3. Binary perceptual study. Participants choose their preferred output between our method and each baseline (LAS (Zheng et al., 2023), SENS (Binninger et al., 2024), and their results post-processed by MeshAnythingV2 (MA) (Chen et al., 2024e)). Our method is preferred by a large margin for both generation and editing.
Operation LAS LAS-MA SENS SENS-MA
Generation 93.3% 94.2% 83.7% 91.0%
Editing 94.4% 96.5% 91.9% 94.4%
Mesh Editing. Based on the results of the perceptual study, together with the sketch metrics reported in Tabs. 2 and 3, our method achieves the best LPIPS and CLIP scores among baselines and reaches the highest rating and preference among users. From the visual comparisons provided in Fig. 7, we notice that baseline methods suffer not only from poor editing results but also from inconsistency in unedited regions. This is because these methods regenerate the whole shape after the edit, can change regions not intended to be edited. Our method solves this problem by decoupling the editing procedure into deletion and addition, ensuring that only edited regions are changed.
Figure 5. Novelty of generated shapes. We retrieve the nearest neighbors of the generated meshes by first aligning them to each train mesh with Iterative Closest Point (ICP) and then computing the Chamfer distance. Our method can generate new shapes different from those in the train set.
Novelty Analysis. Fig. 5 shows a qualitative novelty analysis, retrieving for our generated meshes the 3 nearest neighbors from the training set, based on Chamfer distance. Our generated shapes differ from their nearest neighbors in both overall shape and triangulation, which demonstrates that our method can generate new shapes.
Figure 6. Comparisons between our interactive generation and editing vs. single-step generation. Left: meshes interactively created using our method, which effectively uses multiple localized generation steps to achieve high complexity. Middle: synthetic sketches generated from the final meshes via Canny Edge detection. Right: meshes generated by applying one pass of the addition network (without iterative editing) on the middle sketches. Our interactive editing pipeline enables creating intricate shapes that can be challenging to generate in one step.
Generating Complex Shapes through Interactive Editing. Our interactive mesh editing enables not only an iterative creation process, but its localized editing focus enables construction of complex 3D shapes by decomposing their generation into a sequence of simpler part generations. As a result, artists can use our method to create a larger variety of complex shapes than the single-step generation, as demonstrated in Fig. 6. We refer to our supplementary video for more interactive modelings in action.
4.5. Ablations
Tab. 4 ablates our vertex-aligned speculator design choices. We conduct ablation studies on three variants of the addition network: without speculator, without vertex-aligned decoding (with the same decoding length of $2$), and without joint training with the OPT.
Joint training provides context-rich hidden states as speculator input. We find that the vanilla speculator training scheme with frozen transformer (Wertheimer et al., 2024) does not work in our mesh generation setting, as the hidden states from our pre-trained transformer do not contain enough contextual information. As shown in Tab. 4, by instead training the speculator jointly with the OPT, our model achieves strong mesh quality performance. In contrast, training the speculator with the OPT frozen results in poor performance. This indicates the OPT learns to encode contexts into the hidden states when jointly trained.
Vertex-alignment preserves mesh quality while achieving comparable speedup as vanilla MLP speculator. Comparing no speculator (w/o speculator) and speculator without vertex alignment (w/o vert-alignment), we can see that the speculator, without aligning to vertex tokens, results in a notable speedup, but at large penalty cost in mesh quality and sketch-to-mesh correspondence. By aligning the speculator to vertex tokens, our method reaches the same level of sketch-to-mesh correspondence as the no speculator version while maintaining the naive speculator’s advantage in generation speed. Additionally, applying the vertex-aligned speculator results in slightly higher metrics in mesh quality. This is likely due to the vertex-aligned speculator reducing the complexity of the linear prediction head. Rather than predicting control tokens and all $x$, $y$, and $z$ coordinate tokens, it is trained to predict only control tokens and $x$ coordinate tokens, which reduces the overall prediction error. Thus, equipped with the vertex-aligned speculator, our method achieves strong performance and fast generation speeds (approximately $26.7$ faces per second considering the average face density of the mesh sequences).
Runtime. Our mesh addition, which typically generates $\approx\nicefrac{{1}}{{5}}$ of a shape, runs in about 1 to 5 seconds (depending on the number of faces added), enabling interactive runtimes for editing.
Limitations. While our method produces interactive editing for 3D meshes, our transformer-based approach is bound by the memory size to a maximum sequence length, resulting in a limited number of mesh triangles ($768$ triangles in our experiments) that can be processed. Although there are methods (Hao et al., 2024; Tang et al., 2024) that extend the limit of the token amount in mesh generation, the upper bound (thousands of triangles) is still far below the requirement of modern computer graphics applications (millions of triangles). Thus, we demonstrate the efficacy of our approach on isolated 3D shapes, but would require paradigms in learned mesh representations in order to generate and edit larger-scale scenes.
Table 4. Ablation study. We ablate our speculator, vertex alignment of our speculator (using a common MLP speculator (Wertheimer et al., 2024) with the same length of $2$), and its joint training with our mesh transformer. The unit of CD is 0.001. T/s represents tokens per second and is measured on NVIDIA A100. Our vertex-aligned speculator accelerates generation by 2.16×2.16$\times$2.16 × without loss in mesh quality. Compared to an MLP speculator without vertex-alignment, which suffers from quality loss, our method runs at a comparable speed.
Quality Metrics Speed
[…]
w/o speculator 7.66 0.1765 95.59 32.59 60.7
w/o vert-alignment 9.00 0.1992 94.43 35.65 138.9
w/o joint training 57.13 0.5134 84.52 211.46 125.3
Ours 6.78 0.1837 96.67 29.67 131.1
5. Conclusion
MeshPad introduces interactive editing into autoregressive mesh generation, by decomposing editing into addition and deletion operations on sequentialized mesh tokens. In contrast to existing works which focus on full shape generation, MeshPad enables fine-grained and localized control over the artistic generation process. To enable interactive feedback, our vertex-aligned speculator notably speeds up mesh face generation, running in just a few seconds in response to user sketch updates. Our sketch-conditioned approach also enables ease of use in creation of complex 3D meshes from novice users. We believe that interactive mesh editing represents a significant step toward democratizing content creation, enabling everyone to create artist-reminiscent 3D meshes.
Acknowledgements.
This work was funded by AUDI AG. Lei Li and Angela Dai were supported by the ERC Starting Grant SpatialSem (101076253) and Matthias Nießner by the ERC Consolidator Grant Gen3D (101171131).